THREE NORMALS ARE DRAWN FROM THE POINTS (14,7) TO THE CURVE y2-16X-8Y=0. FIND THE COORDINATES OF FEET OF NORMALS.


 


NOTE. IN EQUNTION 2 IS POWER OF y

Grade:12

2 Answers

Aman Bansal
13 years ago

dear manoj,

let the coordinates of foot of perprndicular be (x1,y1)

differentiating the equation to find the slope of the normal,

=> dy/dx = 8/(y-4)

slope = -(dx/dy)

hence we can find the equation of normal in terms of x1 and y1.

then passing through the (x1,y1)

thus we find a cubic equation on solving which we get,

y1= 0, 16 , -4

thus corresponding x1 = 0, 8 , 3

Don’t forget to approve the answer if it is beneficial to you… !!

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

AMAN BANSAL

 

Ashhad Quadri
8 years ago
  1. Y=mx-2am-am3
As we know that this equation implies for standard  parabola 
But the given equation  is in shifted form 
After simplifying  the equation 
(Y-4 )2=16 (X+1)
Replacing y-4 =Y
And x+1=1
Placing in equation of normal and x=14 and y=7 we get m in cubic find the roots  
Remember point of contact also changes.